# references.bib

### v1

@@ Abstract
Rare diseases affect an estimated 300-400 million people worldwide, yet individual conditions often remain poorly characterized and difficult to diagnose due to their low prevalence and limited clinician familiarity. While computational phenotyping algorithms show promise for automating rare disease detection, their development is hindered by the scarcity of labeled data and biases in existing label sources. Gold-standard labels from registries and expert chart reviews are highly accurate but constrained by selection bias and the cost of manual review. In contrast, labels derived from electronic health records (EHRs) cover a broader range of patients but can introduce substantial noise. To address these challenges, we propose a weakly supervised, transformer-based framework that combines a small set of gold-standard labels with a large volume of iteratively updated silver-standard labels derived from EHR data. This hybrid approach enables the training of a highly accurate and generalizable phenotyping model that scales rare disease detection beyond the scope of individual clinical expertise. Our method is initialized by learning embeddings of medical concepts based on their semantic meaning or co-occurrence patterns in EHRs, which are then refined and aggregated into patient-level representations via a multi-layer transformer architecture. Using two rare pulmonary diseases as a case study, we validate our model on EHR data from Boston Children’s Hospital. Our framework demonstrates notable improvements in phenotype classification, identification of clinically meaningful subphenotypes through patient clustering, and prediction of disease progression compared to baseline methods. These results highlight the potential of our approach to enable scalable identification and stratification of rare disease patients for clinical care and research applications.
@@ 1.   Introduction
Rare diseases are broadly defined as conditions affecting fewer than 1 in 2,000 people in any World Health Organization region or, in the United States, as those affecting fewer than 200,000 people [health2024landscape, wang2024operational]. While individually uncommon, rare diseases collectively impose a significant public health burden, affecting an estimated 300-400 million people worldwide [health2024landscape, marwaha2022guide]. Among these, approximately 30 million people in the United States – 10% of Americans – are living with a rare disease, a prevalence comparable to that of type 2 diabetes [marwaha2022guide, boulanger2020establishing].
@@
Despite their widespread impact, rare diseases – spanning more than 7,000 distinct conditions – remain disproportionately difficult to diagnose. Clinicians may encounter some of these conditions only once, if ever, in their careers, limiting their familiarity with the full spectrum of clinical presentations [mak2024computer, rubinstein2020case]. These challenges contribute to the phenomenon of the “diagnostic odyssey” – a years-long journey marked by inconclusive tests, repeated specialist referrals, and frequent misdiagnoses – experienced by most rare disease patients [bauskis2022diagnostic]. On average, these patients consult between three and ten different physicians, with delays ranging from four to seven years before receiving a correct diagnosis [stoller2018challenge, health2024landscape, mak2024computer]. Prolonged diagnostic delays leave patients without effective treatment for extended periods, increasing the risk of preventable complications, disease progression, and premature mortality [sreih2021diagnostic, gunne2020retrospective, mazzucato2023estimating]. This burden is particularly acute in pediatric populations, as 70% of rare diseases have childhood onset, and 30% of affected children die before the age of five – underscoring the urgent need for earlier and more accurate diagnosis to improve pediatric outcomes and support sustained health and quality of life across the lifespan [eclinicalmedicine2023rare, health2024landscape].
@@
The challenges of rare disease diagnosis are further magnified in the context of rare pulmonary diseases, which remain notoriously difficult to identify due to their symptomatic overlap with more common respiratory conditions. Up to one-third of individuals initially diagnosed with asthma are later found to have been misdiagnosed, with their symptoms more accurately attributed to less prevalent comorbid conditions [gherasim2018confounders, kavanagh2019over]. Pulmonary hypertension (PH), a progressive disorder characterized by high blood pressure in the pulmonary arteries, frequently presents with nonspecific symptoms such as breathlessness, fatigue, and weakness – clinical features that closely resemble those of asthma [ruopp2022diagnosis, galie20162015]. This overlap often leads to delayed diagnosis, by which point irreversible vascular damage may already be present [brown2011delay]. Severe asthma, a distinct and high-burden condition requiring treatment with high-dose inhaled corticosteroids plus a second controller, presents a similarly complex diagnostic challenge [chung2018diagnosis]. Despite accounting for more than one-third of asthma-related deaths, severe asthma remains under-recognized, and its clinical heterogeneity further complicates both timely diagnosis and effective management [levy2014asthma]. Together, these diagnostic pitfalls highlight the limitations of relying solely on clinical expertise and underscore the need for data-driven approaches that can detect subtle, multi-dimensional patterns often missed in routine practice.
@@
Efforts to compile rare disease cases into condition-specific registries have helped consolidate data for research [d2016creating], but these registries are typically too small and narrow in scope to support comprehensive rare disease studies [gliklich2014registries, hageman2023systematic]. Furthermore, because registry inclusion requires a confirmed diagnosis, individuals with atypical presentations or missed diagnoses – who are critical to building more representative datasets for research – are systematically excluded. The widespread adoption of electronic health records (EHRs) has enabled large-scale investigation of rare diseases across diverse patient populations and care settings, capturing a broader spectrum of clinical presentations than traditional registries. EHRs contain rich longitudinal data in both structured (e.g., diagnosis, medication, and procedure codes) and unstructured (e.g., free-text clinical notes from which we extract concept unique identifiers [CUIs]) formats, collectively documenting a patient’s diagnostic journey – including the misdiagnoses and testing patterns that can help characterize rare diseases [garcelon2020electronic]. With access to these data, machine-assisted diagnostic approaches are becoming increasingly integrated into both clinical care and research workflows, with notable success in pulmonary medicine, where they support real-time interpretation of imaging studies [walsh2018deep, huang2019deep], flag high-risk individuals for specialist referrals [kaplan2021artificial], and retrospectively identify undiagnosed patients for inclusion in disease registries and observational studies [geva2017computable].
@@
Central to these efforts is computational phenotyping, which aims to automate the identification of patterns in EHR data that distinguish patients with a given disease (phenotyping) and further stratify them based on clinically meaningful subgroups such as prognosis or treatment response (subphenotyping). Traditionally, phenotyping has relied on rule-based algorithms that apply predefined logical criteria – such as the presence of a specific diagnostic code or concept, use of a relevant medication, or an abnormal lab value – to infer disease status from a patient’s clinical profile [shivade2014review, alzoubi2019review]. While effective for well-characterized diseases with standardized coding practices, these methods translate poorly to rare diseases, which are often clinically heterogeneous and lack codified diagnostic criteria [yang2023machine, banda2018advances].
@@
To overcome the limitations of rule-based approaches, machine learning (ML) and, in particular, deep learning (DL) have emerged as powerful alternatives to enable phenotyping at scale [yang2023machine, callahan2023characterizing]. A key innovation in this space is representation learning, which transforms high-dimensional clinical data into lower-dimensional vector spaces that preserve semantic and contextual relationships [choi2016multi]. Within this framework, medical concepts from both structured and unstructured EHR data are mapped to embeddings pre-trained on co-occurrence patterns and semantic context [weng2019representation]. These concept-level embeddings can be further aggregated into patient-level representations that support downstream phenotyping and subphenotyping tasks [fridgeirsson2023attention].
@@
Despite their success in modeling common diseases, current ML and DL approaches often fail to generalize to rare disease settings. This is due to challenges inherent to both the data and the modeling paradigm. First, EHR data are high-dimensional, sparse, and noisy [banerjee2023machine, schaefer2020use]. Crucially, the presence of a diagnostic code or concept in the EHR does not always indicate a confirmed physician diagnosis – it may be entered for billing purposes, reflect rule-out or provisional conditions, or persist from outdated assessments [yang2024noise, wu2020statistics]. Documentation practices vary widely across providers and institutions, further compounding label inconsistency. These factors degrade the reliability of EHR-derived labels and introduce substantial noise into downstream phenotyping tasks. Second, the dominant ML/DL paradigm for phenotyping is supervised learning, which depends on large volumes of high-quality labeled data – a resource rarely available in rare disease contexts. Supervised models often overfit small, gold-standard cohorts and struggle to generalize to broader, more diverse patient populations. As a result, their clinical utility remains limited. These constraints have fueled growing interest in weakly supervised learning, an approach that leverages large quantities of noisy or partially labeled data to improve model performance in low-label settings.
@@
Among emerging deep learning architectures, transformers have shown particular promise for EHR-based phenotyping due to their ability to model complex temporal dependencies and long-range relationships across irregular clinical events [vaswani2017attention, yang2023transformehr]. Recent models such as BEHRT [li2020behrt], Med-BERT [rasmy2021med], RatchetEHR [hirszowicz2024icu], and Foresight [kraljevic2024foresight] have demonstrated impressive performance on a variety of predictive and classification tasks across both structured and unstructured clinical data. These advances underscore the potential of transformer-based models to capture nuanced disease patterns and support scalable phenotyping. However, despite their promise, transformer models remain largely constrained by the same supervised learning paradigm that limits broader ML/DL applications in rare disease settings. Most existing transformer-based approaches depend heavily on large volumes of clean, labeled training data. As a result, their applicability to low-label, high-noise scenarios remains limited, and their potential for rare disease detection is yet to be fully realized.
@@
To address this gap, we propose a weakly supervised transformer framework for rare disease phenotyping and subphenotyping. Our method combines a small set of expert-validated gold-standard labels with a much larger set of iteratively refined silver-standard labels derived from real-world EHR data. By integrating precise supervision with abundant but noisy signals, the model learns robust patient representations optimized for downstream tasks such as classification and clustering. Using PH and severe asthma as motivating case studies, we demonstrate that our approach improves phenotype detection and enables clinically meaningful subgroup discovery in real-world, data-limited settings.
@@ 2.   Methods
Our end-to-end pipeline integrates representation learning with weak supervision and iterative label refinement to enable accurate and scalable EHR phenotyping. We first identify a high-risk patient cohort and assign initial phenotypic labels using gold- or silver-standard sources (§2.1). Each patient’s longitudinal clinical history is then transformed into a structured input sequence through a multi-step preprocessing pipeline that includes event aggregation, feature selection, and frequency encoding (§2.2). These inputs are processed by a multi-layer transformer encoder that models dependencies among clinical concepts (§2.3). We then aggregate concept-level embeddings to generate patient-level representations, apply a classification head, and iteratively refine the silver-standard labels through weak supervision (§2.4). The framework outputs both a patient-level phenotype prediction and a low-dimensional embedding suitable for clustering and visualization. An overview of the pipeline is shown in Figure 1.
Figure 1: Overview of the proposed weakly supervised transformer framework for EHR-based phenotyping.
@@ 2.1.   Cohort Identification and Labeling
Let $i=1,\ldots,N$ index all patients in the high-risk cohort. Each patient $i$ is assigned a label $y_{i}$ reflecting their phenotype status. Based on the source and reliability of the label, patients are stratified into one of two subcohorts:
@@
    Gold-Standard Cohort: Patients whose disease status has been confirmed through expert physician chart review or inclusion in a dedicated disease registry. These patients are assigned gold-standard labels, denoted $y_{i}^{\text{gold}}$, which serve as high-fidelity references for model training and evaluation.
@@
    Silver-Standard Cohort: Patients with possible but unconfirmed diagnoses. These patients are assigned silver-standard labels, denoted $y_{i}^{\text{silver}}$, inferred from the EHR data. Silver-standard labels can be defined using rule-based heuristics – such as exceeding a threshold number of occurrences of $c^{*}$ – or derived from the probabilistic predictions of automated phenotyping algorithms such as PheNorm [phenorm] or KOMAP [komap]. While these criteria expand the size of the labeled dataset, silver-standard labels are inherently noisier and require iterative refinement.
@@
$y_{i}=\begin{cases}y_{i}^{\text{gold}},&\text{if patient }i\text{ is in the gold-standard cohort},\\ y_{i}^{\text{silver}},&\text{if patient }i\text{ is in the silver-standard cohort}.\end{cases}$
@@ 2.2.3 Feature Engineering and Embedding Construction
Let $\mathcal{C}=\{(c_{1},n_{1}),(c_{2},n_{2}),...,(c_{K},n_{K})\}$ denote the set of unique concepts and their cumulative counts across a patient’s selected time period, whether from $\mathcal{P}$ or $\mathcal{P}^{\prime}$. Each concept $c_{k}\in\mathcal{C}$ is mapped to a vector representation $\mathbf{e}_{k}$ using a pre-trained embedding model (PEM) such as SapBERT [liu2020sapbert], CODER [yuan2022coder], MUGS [li2024multi], or ONCE [komap]:
$\mathbf{e}_{k}=\text{PEM}(c_{k}),\quad\mathbf{e}_{k}\in\mathbb{R}^{d_{\text{input}}}.$ (1)
@@
$\mathbf{e}_{k}^{\text{proj}}=\mathbf{W}^{\text{proj}}\mathbf{e}_{k}+\mathbf{b}^{\text{proj}},\quad\mathbf{e}_{k}^{\text{proj}}\in\mathbb{R}^{d_{\text{model}}},$ (2)
@@
Given the potentially large number of unique concepts in $\mathcal{C}$, we perform feature selection to retain only those most relevant to the target condition. This serves two purposes: (1) reducing noise from unrelated concepts, and (2) lowering computational burden, since transformer attention scales quadratically with the number of input tokens.
To identify relevant features, we compute the cosine similarity between the embedding of each concept and that of the target concept $c^{*}$ for each patient:
@@
$\mathcal{C}^{*}=\{(c_{1},n_{1}),(c_{2},n_{2}),...,(c_{K^{*}},n_{K^{*}})\},\quad\text{where }S(c_{1},c^{*})\geq S(c_{2},c^{*})\geq...\geq S(c_{K^{*}},c^{*}).$
@@
At this stage, we have constructed an aggregated set $\mathcal{C}^{*}$ comprising unique clinical concepts and their corresponding cumulative frequencies, which summarize a patient’s longitudinal medical history. To encode concept frequency – used here as a proxy for clinical significance, capturing aspects such as chronicity or ongoing management – we introduce a frequency-based embedding mechanism.
Each cumulative count $n_{k}$ for concept $c_{k}$ is projected into the model’s embedding space through a two-layer feedforward network with a SwiGLU activation function:
$\mathbf{p}_{k}=\mathbf{W}^{\text{pos}}_{2}\text{SwiGLU}\left(n_{k}\mathbf{W}^{\text{pos}}_{1}+\mathbf{b}^{\text{pos}}_{1}\right)+\mathbf{b}^{\text{pos}}_{2},\quad\mathbf{p}_{k}\in\mathbb{R}^{d_{\text{model}}},$ (4)
@@
$\mathbf{W}^{\text{pos}}_{1}\in\mathbb{R}^{\frac{d_{\text{model}}}{2}\times 1},\quad\mathbf{W}^{\text{pos}}_{2}\in\mathbb{R}^{d_{\text{model}}\times\frac{d_{\text{model}}}{2}},\quad\mathbf{b}^{\text{pos}}_{1}\in\mathbb{R}^{\frac{d_{\text{model}}}{2}},\quad\mathbf{b}^{\text{pos}}_{2}\in\mathbb{R}^{d_{\text{model}}}.$
Unlike traditional positional encodings used in NLP, this representation is grounded in concept frequency rather than token order, offering a tailored signal for clinical models sensitive to the recurrence and persistence of medical events.
@@
The final representation of each selected concept is obtained by summing its embedding and frequency encoding:
$\mathbf{z}_{k}=\mathbf{e}_{k}^{\text{proj}}+\mathbf{p}_{k},\quad\mathbf{z}_{k}\in\mathbb{R}^{d_{\text{model}}}.$ (5)
Here, $\mathbf{z}_{k}$ is the input token for concept $c_{k}$ to the transformer. This formulation allows the model to simultaneously capture semantic similarity across medical concepts and their implicit clinical significance based on frequency. The final patient sequence is:
$\mathbf{Z}_{i}=\{\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{K^{*}}\}.$

### v2

# A Weakly Supervised Transformer for Rare Disease Diagnosis and Subphenotyping from EHRs with Pulmonary Case Studies

@@ Abstract
Rare diseases affect an estimated 300-400 million people worldwide, yet individual conditions remain underdiagnosed and poorly characterized due to their low prevalence and limited clinician familiarity. Computational phenotyping offers a scalable approach to improving rare disease detection, but algorithm development is hindered by the scarcity of high-quality labeled data for training. Expert-labeled datasets from chart reviews and registries are clinically accurate but limited in scope and availability, whereas labels derived from electronic health records (EHRs) provide broader coverage but are often noisy or incomplete. To address these challenges, we propose WEST (WEakly Supervised Transformer for rare disease phenotyping and subphenotyping from EHRs), a framework that combines routinely collected EHR data with a limited set of expert-validated cases and controls to enable large-scale phenotyping. At its core, WEST employs a weakly supervised transformer model trained on extensive probabilistic silver-standard labels – derived from both structured and unstructured EHR features – that are iteratively refined during training to improve model calibration. We evaluate WEST on two rare pulmonary diseases using EHR data from Boston Children’s Hospital and show that it outperforms existing methods in phenotype classification, identification of clinically meaningful subphenotypes, and prediction of disease progression. By reducing reliance on manual annotation, WEST enables data-efficient rare disease phenotyping that improves cohort definition, supports earlier and more accurate diagnosis, and accelerates data-driven discovery for the rare disease community.
Keywords: Weakly Supervised Learning, Transformer Models, Rare Disease, Computational Phenotyping, Subphenotyping, Electronic Health Records
@@ 1.  Introduction
Rare diseases are broadly defined as conditions affecting fewer than 1 in 2,000 people in any World Health Organization region or, in the United States, as those affecting fewer than 200,000 people [1, 2]. While individually uncommon, rare diseases collectively impose a substantial public health burden, affecting an estimated 300-400 million people worldwide [1, 3]. In the United States alone, approximately 30 million individuals – 10% of the population – are living with a rare disease, a prevalence comparable to that of type 2 diabetes [3, 4].
@@
Despite this widespread impact, rare diseases – spanning more than 7,000 distinct conditions – remain disproportionately difficult to diagnose. Many clinicians encounter some of these conditions only once, if ever, in their careers, limiting familiarity with the full spectrum of clinical presentations [5, 6]. These challenges contribute to the so-called “diagnostic odyssey” – a years-long process marked by inconclusive tests, repeated specialist referrals, and frequent misdiagnoses – that most rare disease patients endure [7]. On average, patients consult between three and ten physicians and wait four to seven years before receiving a correct diagnosis [8, 1, 5]. Such delays prevent patients from receiving timely treatment, increase the risk of preventable complications, and contribute to premature mortality [9, 10, 11]. The burden is particularly acute in pediatrics, as 70% of rare diseases present in childhood and 30% of affected children die before age five [12, 1]. There is therefore an urgent need for earlier, more accurate diagnosis to improve outcomes and quality of life across the lifespan.
@@
Diagnostic challenges are even more pronounced in rare pulmonary diseases, which are notoriously difficult to identify due to symptomatic overlap with common respiratory conditions. Up to one-third of individuals initially diagnosed with asthma are later found to have been misdiagnosed, with their symptoms attributable instead to less prevalent comorbid conditions [13, 14]. Pulmonary hypertension (PH), a progressive disorder characterized by elevated pulmonary arterial pressure, frequently presents with nonspecific symptoms such as breathlessness, fatigue, and weakness – features that closely resemble asthma [15, 16]. This overlap often delays recognition until irreversible vascular damage has occurred [17]. Severe asthma, a distinct and high-burden phenotype requiring high-dose inhaled corticosteroids plus a second controller, presents a similarly complex diagnostic challenge [18]. Despite accounting for more than one-third of asthma-related deaths, severe asthma remains under-recognized, and its clinical heterogeneity further complicates timely diagnosis and effective management [19]. Together, these pitfalls highlight the limitations of relying solely on clinical expertise and emphasize the need for data-driven approaches capable of detecting subtle, multi-dimensional patterns often missed in routine practice.
@@
Efforts to consolidate rare disease cases into condition-specific registries have provided valuable research resources [20], but registries are often too small and narrow in scope to support large-scale clinical studies [21, 22]. Moreover, because inclusion requires a confirmed diagnosis, patients with atypical presentations or missed diagnoses – those most crucial for building representative datasets – are systematically excluded. The widespread adoption of electronic health records (EHRs) has enabled rare disease investigation at scale, capturing a broader spectrum of clinical presentations than traditional registries. EHRs contain rich longitudinal data in both structured (e.g., diagnosis, medication, and procedure codes) and unstructured (e.g., free-text notes from which concept unique identifiers [CUIs] can be extracted) formats, documenting a patient’s diagnostic journey – including misdiagnoses and testing patterns that can illuminate rare disease phenotypes [23]. Leveraging these data, machine-assisted diagnostic approaches are increasingly integrated into research and clinical workflows, with notable success in pulmonary medicine: they support real-time imaging interpretation [24, 25], flag high-risk patients for specialist referrals [26], and retrospectively identify undiagnosed individuals for registries and observational studies [27].
@@
Central to such efforts is computational phenotyping, which seeks to automate the identification of disease patterns in EHR data and further stratify patients into clinically meaningful subgroups based on prognosis or treatment response. Traditionally, phenotyping has relied on rule-based algorithms that apply predefined logical criteria – such as diagnostic codes, relevant medications, or abnormal lab values – to infer disease status [28, 29]. While effective for well-characterized diseases with standardized coding, these methods perform poorly in rare diseases, which are clinically heterogeneous and often lack codified diagnostic criteria [30, 31].
@@
To address these limitations, machine learning (ML) and especially deep learning (DL) have emerged as powerful alternatives for large-scale phenotyping [30, 32]. A key advance in this domain is representation learning, which transforms high-dimensional clinical data into lower-dimensional embeddings that preserve semantic and contextual relationships [33]. Within this framework, medical concepts from both structured and unstructured data are mapped to embeddings pre-trained on co-occurrence and semantic context [34]. These concept-level embeddings can then be aggregated into patient-level representations to support downstream phenotyping and subphenotyping tasks [35].
@@
Despite their success in modeling common diseases, ML and DL methods often struggle in rare disease contexts due to both data and modeling challenges. EHR data are inherently high-dimensional, sparse, and noisy [36, 37]. Importantly, the presence of a diagnostic code or concept in the EHR does not guarantee a confirmed diagnosis: codes may be entered for billing, used provisionally, or persist from outdated assessments [38, 39]. For rare diseases, which are inconsistently documented and frequently underrecognized, these issues are particularly acute. Variability in documentation practices across providers and institutions further complicates the curation of accurate large-scale training sets. Together, these factors introduce substantial noise into downstream phenotyping. Compounding these challenges, most ML/DL methods for phenotyping rely on supervised learning, which requires large, high-quality labeled datasets – a resource rarely available in rare disease research [40]. Models trained on small gold-standard cohorts often overfit and fail to generalize, limiting clinical utility. These constraints have fueled growing interest in weakly supervised learning, which exploits large collections of noisy or partially labeled data to enhance model robustness under low-label conditions.
@@
Among emerging DL architectures, transformers have shown particular promise for EHR phenotyping due to their ability to capture complex temporal dependencies and long-range relationships across irregular clinical events [41, 42]. Models such as BEHRT [43], Med-BERT [44], RatchetEHR [45], and Foresight [46] have achieved strong performance across predictive and classification tasks in both structured and unstructured data. These advances underscore the potential of transformer-based models for scalable phenotyping. Yet, despite their promise, transformer models remain constrained by the same supervised learning paradigm that limits other ML/DL applications in rare diseases. Existing approaches still require large volumes of clean, labeled data, limiting applicability in low-label, high-noise scenarios. Their potential for rare disease detection is therefore far from fully realized.
@@
To bridge this gap, we propose a weakly supervised transformer (WEST) framework for rare disease phenotyping and subphenotyping from EHRs. The method combines a small set of expert-validated gold-standard labels with a much larger set of iteratively refined silver-standard labels derived from real-world EHR data. By integrating precise supervision with abundant but noisy signals, WEST learns robust patient representations optimized for downstream classification and clustering tasks. Importantly, the framework performs effectively even when only positive gold-standard cases are available, making it applicable in settings where registries exist but large-scale chart review is infeasible. Using PH and severe asthma as motivating case studies, we demonstrate that this approach improves phenotype detection and enables clinically meaningful subgroup discovery in real-world, data-limited settings.
@@ 2.  Methods
Our end-to-end WEST phenotyping pipeline integrates representation learning with weak supervision and iterative label refinement to enable data-efficient EHR phenotyping. We first identify a high-risk patient cohort and assign initial phenotypic labels using gold- or silver-standard sources (§2.1). Each patient’s longitudinal clinical history is then transformed into a structured input sequence through a multi-step pre-processing pipeline that includes event aggregation, feature selection, and frequency encoding (§2.2). These inputs are processed by a multi-layer transformer encoder that models dependencies among clinical concepts (§2.3). We then aggregate concept-level embeddings to generate patient-level representations, apply a classification head, and iteratively refine the silver-standard labels through weak supervision (§2.4). The framework outputs both a patient-level phenotype prediction and a low-dimensional embedding suitable for clustering and visualization. An overview of the pipeline is shown in Figure 1.
Figure 1: Overview of the WEST phenotyping pipeline.
@@ 2.1.  Cohort Identification and Labeling
Let $i=1,\ldots,N$ index all patients in the high-risk cohort. Each patient $i$ is assigned a label $y_{i}$ reflecting their phenotype status. Based on the source and reliability of the label, patients are stratified into two cohorts:
@@
    Gold-Standard Cohort: Patients whose disease status has been confirmed through expert physician chart review or inclusion in a dedicated disease registry. These patients are assigned gold-standard labels, denoted $y_{i}^{\text{gold}}$, which serve as high-fidelity references for model training and evaluation. We allow this set to be small to ensure that the WEST pipeline is label efficient.
@@
    Silver-Standard Cohort: Patients with possible but unconfirmed diagnoses. These patients are assigned silver-standard labels, denoted $y_{i}^{\text{silver}}$, inferred from the EHR data. Silver-standard labels can be defined using rule-based heuristics – such as exceeding a threshold number of occurrences of $c^{*}$ – or derived from the probabilistic predictions of unsupervised automated phenotyping algorithms such as PheNorm [47] or KOMAP [48]. While these criteria expand the size of the labeled dataset, silver-standard labels are inherently noisier and require iterative refinement.
@@
[…]
@@ 2.2.3 Feature Engineering and Embedding Construction
Let $\mathcal{C}=\{(c_{1},n_{1}),(c_{2},n_{2}),...,(c_{K},n_{K})\}$ denote the set of unique concepts and their cumulative counts across a patient’s selected time period, whether from $\mathcal{P}$ or $\mathcal{P}^{\prime}$. Each concept $c_{k}\in\mathcal{C}$ is mapped to a vector representation $\mathbf{e}_{k}$ using a pre-trained embedding model (PEM) for clinical concepts such as SapBERT [49], CODER [50], MUGS [51], or ONCE [48]:
$\mathbf{e}_{k}=\text{PEM}(c_{k}),\qquad\mathbf{e}_{k}\in\mathbb{R}^{d_{\text{input}}}.$ (1)
@@
$\mathbf{e}_{k}^{\text{proj}}=\mathbf{W}^{\text{proj}}\mathbf{e}_{k}+\mathbf{b}^{\text{proj}},\qquad\mathbf{e}_{k}^{\text{proj}}\in\mathbb{R}^{d_{\text{model}}},$ (2)
@@
Given the potentially large number of unique concepts in $\mathcal{C}$, we perform feature selection to retain only those most relevant to the target condition. This serves two purposes: (1) reducing noise from unrelated concepts, and (2) lowering computational burden, since transformer attention scales quadratically with the number of input tokens. To identify relevant features, we compute the cosine similarity between the embedding of each concept and that of the target concept $c^{*}$, representing the disease condition of interest:
@@
$\mathcal{C}^{*}=\{(c_{1},n_{1}),(c_{2},n_{2}),...,(c_{K^{*}},n_{K^{*}})\},\qquad\text{where }S(c_{1},c^{*})\geq S(c_{2},c^{*})\geq...\geq S(c_{K^{*}},c^{*}).$
@@
At this stage, we have constructed an aggregated set $\mathcal{C}^{*}$ comprising unique clinical concepts and their corresponding cumulative frequencies, which summarize a patient’s longitudinal medical history. To encode concept frequency – serving as a proxy for clinical significance, capturing aspects such as chronicity or ongoing management – we introduce a frequency-based embedding mechanism.
Each patient-specific cumulative count $n_{ki}$ for concept $c_{k}$ is projected into the model’s embedding space through a two-layer feedforward network with a SwiGLU activation function, a gated variant of the linear unit shown to improve expressivity and training stability in transformer feedforward layers [52]:
$\mathbf{p}_{ki}=\mathbf{W}^{\text{pos}}_{2}\text{SwiGLU}\left(n_{ki}\mathbf{W}^{\text{pos}}_{1}+\mathbf{b}^{\text{pos}}_{1}\right)+\mathbf{b}^{\text{pos}}_{2},\qquad\mathbf{p}_{ki}\in\mathbb{R}^{d_{\text{model}}},$ (4)
@@
$\mathbf{W}^{\text{pos}}_{1}\in\mathbb{R}^{\frac{d_{\text{model}}}{2}\times 1},\qquad\mathbf{W}^{\text{pos}}_{2}\in\mathbb{R}^{d_{\text{model}}\times\frac{d_{\text{model}}}{2}},\qquad\mathbf{b}^{\text{pos}}_{1}\in\mathbb{R}^{\frac{d_{\text{model}}}{2}},\qquad\mathbf{b}^{\text{pos}}_{2}\in\mathbb{R}^{d_{\text{model}}}.$
Unlike traditional positional encodings used in natural language processing (NLP), this representation is grounded in concept frequency rather than token order, offering a tailored signal for clinical models sensitive to the recurrence and persistence of medical events.
@@
The final representation of each selected concept is obtained by summing its embedding and patient-specific frequency encoding:
$\mathbf{z}_{ki}=\mathbf{e}_{k}^{\text{proj}}+\mathbf{p}_{ki},\qquad\mathbf{z}_{ki}\in\mathbb{R}^{d_{\text{model}}}.$ (5)
Here, $\mathbf{z}_{ki}$ is the input token for concept $c_{k}$ for patient $i$ to the transformer. If concept $c_{k}$ is not observed for patient $i$, we set $\mathbf{z}_{ki}=0$. This formulation allows the model to simultaneously capture semantic similarity across medical concepts and their implicit clinical significance based on frequency. The final patient sequence is:
$\mathbf{Z}_{i}=\{\mathbf{z}_{1i},\mathbf{z}_{2i},...,\mathbf{z}_{K^{*}i}\}.$